Marvel Parts, Inc., manufactures auto accessories. One of the companys products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 2,850 hours each month to produce 1,900 sets of covers. The standard costs associated with this level of production are: Per Set of Covers Total Direct materials Direct labor Variable manufacturing overhead (based on direct $ 42,560 $22.40 s 51,300 27.00 $ 6,8403.60 $53.00 labor-hours) During August, the factory worked only 2,800 direct labor-hours and produced 2,000 sets of covers. The following actual costs were recorded during the month: Per Set Total of Covers Direct materials (12,000 yards) Direct labor Variable manufacturing overhead S 45,600 $22.80 S 49,000 24.50 $ 7,000 3.50 $50.80 At standard, each set of covers should require 5.6 yards of material. All of the materials purchased during the month were used in production. Required: 1. Compute the materials price and quantity variances for August. 2. Compute the labor rate and efficiency variances for August. 3. Compute the variable overhead rate and efficiency variances for August. (Indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance). Input all amounts as positive values.)
1 Materials price variance Materials quantity variance 2 Labor rate variance Labor efficiency variance 3 Variable overhead rate variance Variable overhead efficiency variance
Answer #1

Ans-1-a- Material Price Variance = Actual quantity*Standard price-Actual quantity* Actual price

= 12,000* $4- 12,000*$3.8

=$48,000-$45,600

=$2,400 (F)

Standard price = Standard material cost/ Standard material quantity

=$42,560/ (1,900*5.6)

=$42,560/ 10,640

=$4 per yard

Actual price= Actual material cost/ Actual material quantity

=$45,600/12,000

=$3.8 per yard

1-b- Material quantity variance= (Standard quantity allowed- Actual quantity)*Standard price

=(5.6*2,000)-12,000)*4

=(11,200-12,000)*4

=$3,200 (U)

2-a- Labor rate variance= Actual hours *(Standard rate- Actual rate)

=2,800*(18-17.5)

=$1,400 (U)

Standard rate= Standard labor cost/ standard labor hours

=$51,300/ 2,850

=$18 per hour

Actual rate= Actual labor cost/ Actual labor hours

=$49,000/ 2,800

=$17.5 per hour

2-b- Labor efficiency variance= Standard rate*(Standard hours - Actual hours)

=18*(3,000-2,800)

=$3,600 (F)

Standard hour per unit= Standard hours/ Standard production

=2,850/1,900

=1.50 per unit

Standard hours= Standard hour per unit * Actual production

=1.50*2,000

=3,000 hours

3-a- Variable overhead rate variance= Actual labor hours *(Standard rate- Actual rate)

=2,800* (2.4-2.5)

=$280 (U)

Standard rate= Standard variable overhead cost/ standard labor hours

=$6,840/2,850

=$2.40 per hour

Actual rate= Actual variable overhead cost. Actual labor hours

=$7,000/2,800

=$2.50 per hour

3-b- Variable overhead efficiency variance= Standard rate* (Standard hours-Actual hours)

=2.40(3,000-2,800)

=$480 (F)

Tabular Representation:-

1 Material price variance $2,400(F)
Material quantity variance $3,200 (U)
2 Labor rate variance $1,400(U)
Labor efficiency variance $3,200 (F)
3 Variable overhead rate variance $280 (U)
Variable overhead efficiency variance $480 (F)